# Existence of cohomogeneity one Einstein metrics

### v4

@@ 1. Introduction
A Riemannian metric $g$ is Einstein if $\mathrm{Ric}(g)=\Lambda g$ for some constant $\Lambda$. The search for Einstein metrics continues to be a challenging problem. A milestone in this field is the existence theory of Kähler–Einstein metrics, with fundamental results including the resolution of the Calabi conjecture in the non-positive case [Aub76, Yau78] and the Yau–Tian–Donaldson conjecture for Fano manifolds [CDS15, Tia15]. While these results have profoundly shaped our understanding of Einstein metrics, many challenges remain in the non-Kähler setting. Significant efforts have been devoted to constructing Einstein metrics in the real Riemannian context. One fruitful approach involves considering homogeneous spaces, where the high degree of symmetry simplifies the Einstein equation to an algebraic system. In the case of homogeneous spaces, a general existence theory has been developed in [WZ86, BWZ04, Böh04]. It is known in [BK06] that each compact simply connected homogeneous space of dimension less than $12$ admits a homogeneous Einstein metric. For low dimensional homogeneous Einstein manifolds up to dimension $7$; see [Jen69, Gor77, ADF96, NR99, NR03, Nik04].
To explore Einstein metrics with large symmetry but beyond homogeneity, a natural direction is cohomogeneity one Einstein metrics, where the isometry group acts with principal orbits $\mathsf{G}/\mathsf{K}$ of codimension one. The cohomogeneity one condition transforms the Einstein equation to a system of ODEs. A landmark example is Page’s metric on $\mathbb{CP}^{2}\sharp\overline{\mathbb{CP}}^{2}$, the firstly known inhomogeneous positive Einstein metric [Pag78]. Page’s metric is generalized in a series of works including [BB82, KS86, PP87, KS88, WW98]. In [Böh98], new cohomogeneity one Einstein metrics were found on some low dimensional manifolds, including spheres and $\mathbb{HP}^{2}\sharp\overline{\mathbb{HP}}^{2}$. The latter space is the quaternionic analogue of Page’s metric. Böhm’s metric is generalized to all high dimensional $\mathbb{HP}^{m+1}\sharp\overline{\mathbb{HP}}^{m+1}$ in [Chi24], confirming numerical evidence in [PP86].
@@
The compact cohomogeneity one manifold $M$ of dimension $n+1=d_{1}+d_{2}+1$ is an $\mathbb{S}^{d_{1}+1}$-bundle over $\mathcal{Q}^{d_{2}}$. Since the manifold $M$ is formed by gluing two copies of a disk bundle over $\mathcal{Q}^{d_{2}}$ along $\mathsf{G}/\mathsf{K}$, it is also referred to as a double disk bundle associated to the group triple $(\mathsf{K},\mathsf{H},\mathsf{G})$. We consider double disk bundles of two summands type, where the cohomogeneity one Einstein equation admits an ansatz consisting of two metric components $f_{1}^{2}$ and $f_{2}^{2}$, respectively for $\mathfrak{h}/\mathfrak{k}$ and $\mathfrak{g}/\mathfrak{h}$. Cases where the isotropy representations of principle orbits split into two irreducible summands are included. The cohomogeneity one Einstein equation on $M$ is characterized by the structural triple $(d_{1},d_{2},A)$ associated with $\mathsf{G}/\mathsf{K}$, where the constant $A\geq 0$ is the norm square of the fundamental tensor in the theory of Riemannian submersion [O’N66]. With such a geometric setting, we prove the following existence theorem.
@@
For any $(d_{1},d_{2})$ with $d_{2}\geq d_{1}\geq 2$, there exists a constant $\chi_{d_{1},d_{2}}\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right]$ such that if $\mathsf{G}/\mathsf{K}$ is a principal orbit with the structural triple $(d_{1},d_{2},A)$ and $A\in(0,\chi_{d_{1},d_{2}})$, there exists at least one cohomogeneity one Einstein metric on $M$.
The case where $d_{1}=1$ was well studied in [BB82] and many aforementioned works that followed, where it was not even assumed that the hypersurface of $M$ is a homogeneous space. We hence consider $d_{1}\geq 2$ in this paper. Note the condition $d_{2}\geq d_{1}$ does not restrict generality. If $d_{2}<d_{1}$, then by slice theorem, a tubular neighborhood around $\mathcal{Q}^{d_{2}}$ is identified as a vector bundle whose rank is larger than $d_{2}$. Therefore, the slice representation of $\mathsf{H}$ contains at least two trivial summands, meaning $\mathsf{H}$ does not act transitively on $\mathsf{H}/\mathsf{K}$. By the assumption of $\mathsf{G}/\mathsf{K}$ being of two summands type, the principal orbit is necessarily a Riemann product $\mathbb{S}^{d_{1}}\times\mathcal{Q}^{d_{2}}$ and $A=0$. Then $M=\mathbb{S}^{d_{1}+1}\times\mathcal{Q}^{d_{2}}$ obviously admits a product metric that is Einstein.
The constant $\chi_{d_{1},d_{2}}$ is determined by minimizing certain rational functions. The explicit formula of $\chi_{d_{1},d_{2}}$ for a general $(d_{1},d_{2})$ is fairly complicated. In an upcoming paper [Chi25b], we derive some estimates of $\chi_{d_{1},d_{2}}$ to provide some concrete new examples of compact Einstein manifolds.
This paper is organized as follows. In Section 2, we introduce the dynamical system for cohomogeneity one Einstein metrics of two summands type. In Section 3, we perform local analysis on critical points representing the initial and terminal conditions. In Section 4, we generalize the global analysis in [Chi24] and prove the main theorem. For the sake of readability, formula and computations obtained by Maple are presented in the Appendix.
@@ 2. The cohomogeneity one Einstein equation
For a group triple $(\mathsf{K},\mathsf{H},\mathsf{G})$ as in (1.1), let $\mathsf{G}$ be compact and semisimple. Let $\tilde{b}$ a normal homogeneous metric on $\mathsf{G}/\mathsf{K}$. If $\mathsf{G}$ is simple, let $\tilde{b}$ be the standard homogeneous metric induced by the negative Killing form $-B_{\mathfrak{g}}$. Assume that there exists a $\tilde{b}$-orthogonal decomposition $\mathfrak{g}=\mathfrak{h}/\mathfrak{k}\oplus\mathfrak{g}/\mathfrak{h}$, where each summand is $K$-invariant (not necessarily irreducible). Then, there exists a 2-parameter family of homogeneous metrics on $\mathsf{G}/\mathsf{K}$ with each parameter controls $\mathbb{S}^{d_{1}}$ and $\mathcal{Q}^{d_{2}}$. We further assume that the Ricci endomorphism of $\mathsf{G}/\mathsf{K}$ is diagonal for such a 2-parameter family. The background metric $b$ used in this paper is rescaled from $\tilde{b}$, so that the homogeneous metric
@@
has the Ricci endomorphism $\mathrm{diag}(r_{1}I_{d_{1}},r_{2}I_{d_{2}})$, where
@@
With a principal orbit $\mathsf{G}/\mathsf{K}$ as above, each $f_{i}$ is a function defined on the $1$ dimensional orbit space $[0,t_{*}]$. From [EW00], the shape operator of $\mathsf{G}/\mathsf{K}$ in $M$ is $L:=\mathrm{diag}\left(\frac{\dot{f}_{1}}{f_{1}}I_{d_{1}},\frac{\dot{f}_{2}}{f_{2}}I_{d_{2}}\right)$. The cohomogeneity one Einstein equation is
@@
on $M$ if the following initial and terminal conditions are satisfied.
(2.7) $\lim\limits_{t\to 0}(f_{1},\dot{f_{1}},f_{2},\dot{f_{2}})=(0,1,f,0),\quad\lim\limits_{t\to t_{*}}(f_{1},\dot{f_{1}},f_{2},\dot{f_{2}})=(0,-1,\bar{f},0),\quad f,\bar{f}>0.$
Follow [Chi24]. Let $d\eta=\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}dt$. Define
$X_{1}:=\frac{\frac{\dot{f_{1}}}{f_{1}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},\quad X_{2}:=\frac{\frac{\dot{f_{2}}}{f_{2}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},\quad Y:=\frac{\frac{1}{f_{1}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},\quad Z:=\frac{\frac{f_{1}}{f_{2}^{2}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}}.$
Define functions
@@
Equations (2.3) are transformed to
(2.9) $\begin{split}X_{i}^{\prime}&=X_{i}H\left(G+\frac{1-H^{2}}{n}-1\right)+R_{i}-\frac{1-H^{2}}{n},\quad i=1,2;\\ Y^{\prime}&=Y\left(H\left(G+\frac{1-H^{2}}{n}\right)-X_{1}\right),\\ Z^{\prime}&=Z\left(H\left(G+\frac{1-H^{2}}{n}\right)+X_{1}-2X_{2}\right).\\ \end{split}$
@@
(2.10) $\mathcal{C}:\frac{d_{1}d_{2}}{n}(X_{1}-X_{2})^{2}+d_{1}R_{1}+d_{2}R_{2}=1-\frac{1}{n}.$
@@
(2.11) $\mathcal{C}:\frac{1}{n-1}\left(G-H^{2}+d_{1}R_{1}+d_{2}R_{2}\right)=\frac{1-H^{2}}{n}.$
By (2.9), we can set $Y$ and $Z$ to be non-negative without loss of generality. Furthermore, since
(2.12) $H^{\prime}=(H^{2}-1)\left(G+\frac{1}{n}(1-H^{2})\right)=(H^{2}-1)\left(\frac{1}{n}+\frac{d_{1}d_{2}}{n}(X_{1}-X_{2})^{2}\right),$
@@
is flow invariant. We consider the dynamical system (2.9) on $\mathcal{E}$. Observe that the subsystem restricted on $\mathcal{E}\cap\{H=1\}$ is the cohomogeneity one Ricci-flat equation with the same group triple. To retrieve the original system, we apply [Chi24, Equation (2.10)].

### v5

@@ 1. Introduction
A Riemannian metric $g$ is Einstein if $\mathrm{Ric}(g)=\Lambda g$ for some constant $\Lambda$. The search for Einstein metrics continues to be a challenging problem. According to the sign of $\Lambda$, one speaks of positive ($\Lambda>0$), Ricci-flat ($\Lambda=0$), or negative ($\Lambda<0$) Einstein metrics.
A milestone in this field is the existence theory of Kähler–Einstein metrics, with fundamental results including the resolution of the Calabi conjecture in the non-positive case [Aub76, Yau78] and the Yau–Tian–Donaldson conjecture for Fano manifolds [CDS15, Tia15]. While these results have profoundly shaped our understanding of Einstein metrics, many challenges remain in the non-Kähler setting. Significant efforts have been devoted to constructing Einstein metrics in the real Riemannian context. One fruitful approach involves considering homogeneous spaces, where the high degree of symmetry simplifies the Einstein equation to an algebraic system. In the case of homogeneous spaces, a general existence theory has been developed in [WZ86, BWZ04, Böh04]. It is shown in [BK06] that each compact simply connected homogeneous space of dimension less than $12$ admits a homogeneous Einstein metric. For low dimensional homogeneous Einstein manifolds up to dimension $7$, see [Jen69, Gor77, ADF96, NR99, NR03, Nik04].
To explore Einstein metrics with large symmetry but beyond homogeneity, a natural direction is cohomogeneity one Einstein metrics, where the isometry group acts with principal orbits $\mathsf{G}/\mathsf{K}$ of codimension one. The cohomogeneity one condition transforms the Einstein equation to a system of ODEs. A landmark example is Page’s metric on $\mathbb{CP}^{2}\sharp\overline{\mathbb{CP}}^{2}$, the first known inhomogeneous positive Einstein metric [Pag78]. Page’s metric was generalized in a series of works including [BB82, KS86, PP87, KS88, WW98]. In [Böh98], new cohomogeneity one Einstein metrics were found on some low dimensional manifolds, including spheres and $\mathbb{HP}^{2}\sharp\overline{\mathbb{HP}}^{2}$. The Einstein metric on the latter space can be viewed as a quaternionic counterpart of Page’s metric. Böhm’s metric is generalized to all high dimensional $\mathbb{HP}^{m+1}\sharp\overline{\mathbb{HP}}^{m+1}$ in [Chi24], confirming numerical evidence in [PP86].
@@
The compact cohomogeneity one manifold $M$ of dimension $n+1=d_{1}+d_{2}+1$ is an $\mathbb{S}^{d_{1}+1}$-bundle over $\mathcal{Q}^{d_{2}}$. Since the manifold $M$ is formed by gluing two copies of a disk bundle over $\mathcal{Q}^{d_{2}}$ along $\mathsf{G}/\mathsf{K}$, it is also referred to as a double disk bundle associated to the group triple $(\mathsf{K},\mathsf{H},\mathsf{G})$. We consider double disk bundles of two summands type, where the cohomogeneity one Einstein equation admits an ansatz consisting of two metric components $f_{1}^{2}$ and $f_{2}^{2}$, respectively for $\mathfrak{h}/\mathfrak{k}$ and $\mathfrak{g}/\mathfrak{h}$. Cases where the isotropy representations of principal orbits split into two irreducible summands are included. The cohomogeneity one Einstein equation on $M$ is characterized by the structural triple $(d_{1},d_{2},A)$ associated with $\mathsf{G}/\mathsf{K}$, where the constant $A\geq 0$ is the norm square of the fundamental tensor in the theory of Riemannian submersions [O’N66]. With such a geometric setting, we prove the following existence theorem.
@@
For any $(d_{1},d_{2})$ with $d_{1},d_{2}\geq 2$, there exists a constant $\chi_{d_{1},d_{2}}\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right]$ such that if $\mathsf{G}/\mathsf{K}$ is a principal orbit with the structural triple $(d_{1},d_{2},A)$ and $A\in(0,\chi_{d_{1},d_{2}})$, there exists at least one cohomogeneity one Einstein metric on $M$.
The constant $\chi_{d_{1},d_{2}}$ is determined by several algebraic functions of $(d_{1},d_{2})$. The explicit formula of $\chi_{d_{1},d_{2}}$ for a general $(d_{1},d_{2})$ is fairly complicated. We present some examples in Appendix C. In an upcoming paper [Chi25a], we derive some estimates of $\chi_{d_{1},d_{2}}$ to provide more examples of new compact Einstein manifolds.
Remark 1.2.
The case $d_{1}=1$ was already extensively studied in the work of Bérard-Bergery [BB82] and subsequent works [KS86, PP87, KS88, WW98], where no homogeneity assumption on the hypersurface was required. Hence, our theorem is stated under the assumption $d_{1}\geq 2$. In principle, our result could be generalized to sphere bundles over an inhomogeneous Einstein manifold. However, in such a setting, the connection on the sphere bundle (as the hypersurface) needs to be Yang–Mills, with the pointwise norm of its curvature being constant; see [Bes07, Theorem 9.73]. How frequently such sphere bundles arise outside the homogeneous fibration context is, to the best of the author’s knowledge, unclear.
Remark 1.3.
By Myers’ Theorem [Mye41], a compact manifold admitting a positive Einstein metric must have a finite fundamental group. If $d_{2}=1$, the base space is necessarily $\mathbb{S}^{1}$, which implies that the fundamental group $\pi_{1}(M)$ is not finite. Hence, such a double disk bundle $M$ cannot admit any metric with positive Ricci curvature, and we therefore exclude the case $d_{2}=1$.
Remark 1.4.
Theorem 1.1 also applies when $A=0$, in which case one simply recovers the product metric on $\mathbb{S}^{d_{1}+1}\times\mathcal{Q}^{d_{2}}$. Therefore, Einstein metrics obtained in Theorem 1.1 with $A>0$ may be regarded as analogues of the product Einstein metric on $\mathbb{S}^{d_{1}+1}\times\mathcal{Q}^{d_{2}}$. It is known that for $5\leq n\leq 8$, there exist infinitely many cohomogeneity one Einstein metrics on $\mathbb{S}^{d_{1}+1}\times\mathcal{Q}^{d_{2}}$ [Böh98]. Whether additional non-standard Einstein metrics exist on higher dimensional product spaces is beyond the scope of this paper.
Böhm’s work [Böh98] established the existence of infinitely many cohomogeneity one Einstein metrics by exploiting a local spiraling mechanism: when certain critical points on the Ricci–flat subsystem are foci, integral curves can wind arbitrarily many times around the cone solution. An intermediate value argument combined with symmetry then produces infinitely many heteroclines that represent complete Einstein metrics on $\mathbb{S}^{d_{1}+1}\times\mathcal{Q}^{d_{2}}$.
In contrast, for the broader class of examples considered here (those with $A>0$), no examples are known where the relevant critical points remain foci, so the local spiraling argument is unavailable. Hence our approach is necessarily different: by constructing global barriers and capturing a distinguished integral curve, together with an intermediate value argument and symmetry, we prove the existence of at least one heterocline that represents a complete Einstein metric on $M$.
This paper is organized as follows. In Section 2, we introduce the dynamical system for cohomogeneity one Einstein metrics of two summands type. In Section 3, we perform local analysis on critical points representing the initial and terminal conditions. In Section 4, we generalize the global analysis in [Chi24] and prove the main theorem. For the sake of readability, formulas and computations obtained by Maple are presented in Appendix B.
@@ 2. The cohomogeneity one Einstein equation
For a group triple $(\mathsf{K},\mathsf{H},\mathsf{G})$ as in (1.1), let $\mathsf{G}$ be compact and semisimple. Let $\tilde{b}$ be a normal homogeneous metric on $\mathsf{G}/\mathsf{K}$, that is, a metric induced from an $\mathrm{Ad}(\mathsf{G})$-invariant inner product on the Lie algebra $\mathfrak{g}$. If $\mathsf{G}$ is simple, we take $\tilde{b}$ to be the standard homogeneous metric induced by the negative Killing form $B_{\mathfrak{g}}$. Assume that there exists a $\tilde{b}-$orthogonal decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{h}/\mathfrak{k}\oplus\mathfrak{g}/\mathfrak{h}$, where the latter two summands are $K$-invariant (not necessarily irreducible). Then there exists a 2-parameter family of homogeneous metrics on $\mathsf{G}/\mathsf{K}$ with the parameters controlling $\mathbb{S}^{d_{1}}$ and $\mathcal{Q}^{d_{2}}$. We further assume that the Ricci endomorphism of $\mathsf{G}/\mathsf{K}$ is diagonal for such a 2-parameter family. We fix a background metric $b$ by rescaling $\tilde{b}$ so that the homogeneous metric
@@
has Ricci endomorphism $\mathrm{diag}(r_{1}I_{d_{1}},r_{2}I_{d_{2}})$, where
@@
With a principal orbit $\mathsf{G}/\mathsf{K}$ as above, each $f_{i}$ is a function defined on the $1$-dimensional orbit space parametrized by $t$. From [EW00], the shape operator of $\mathsf{G}/\mathsf{K}$ in $M$ is $L:=\mathrm{diag}\left(\frac{\dot{f}_{1}}{f_{1}}I_{d_{1}},\frac{\dot{f}_{2}}{f_{2}}I_{d_{2}}\right)$. The cohomogeneity one Einstein equation is
@@
on $M$ if $\mathsf{G}/\mathsf{K}$ collapses to $\mathsf{G}/\mathsf{H}$ on both ends. By [EW00] and more recently [VZ22], the smooth extension to the singular orbit requires the following initial and terminal conditions.
(2.7) $\lim\limits_{t\to 0}(f_{1},\dot{f_{1}},f_{2},\dot{f_{2}})=(0,1,\underline{f},0),\quad\lim\limits_{t\to t_{*}}(f_{1},\dot{f_{1}},f_{2},\dot{f_{2}})=(0,-1,\overline{f},0),\quad\underline{f},\overline{f}>0.$
For a complete solution defined on $0,t_{*}]$, many geometric quantities, including the mean curvature $\mathrm{tr}{(L)}$, diverge at both ends of the interval. To compactify the phase space of the dynamical system, we introduce a natural parameter $\eta$ by
$d\eta=\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}dt.$
Because $\mathrm{tr}(L)$ blows up as $t\to 0$ and $t\to t_{*}$, the interval $0,t_{*}]$ is mapped to the entire real line in the $\eta$–coordinate. Under this change of variable, equations (2.3)–(2.4) are transformed into a dynamical system on a compact phase space, which is well suited for the subsequent global analysis. Specifically, define
$X_{1}:=\frac{\frac{\dot{f_{1}}}{f_{1}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},\quad X_{2}:=\frac{\frac{\dot{f_{2}}}{f_{2}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},$
$Y:=\frac{\frac{1}{f_{1}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}},\quad Z:=\frac{\frac{f_{1}}{f_{2}^{2}}}{\sqrt{\mathrm{tr}^{2}(L)+n\Lambda}}.$
Each $X_{i}$ is the normalized principal curvature, while the metric components are encoded in $Y$ and $Z$. Define functions
@@
Up to the normalization factor, the function $G$ represents the squared norm of the shape operator, while $H$ corresponds to the mean curvature. The homogeneous Ricci curvatures of the principal orbit along the fiber and base directions are respectively captured by $R_{1}$ and $R_{2}$.
Here and in what follows, the prime notation () denotes differentiation with respect to $\eta$. Equations (2.3) are transformed to
[…]
@@
(2.10) $\mathcal{C}:=\left\{\frac{d_{1}d_{2}}{n}(X_{1}-X_{2})^{2}+d_{1}R_{1}+d_{2}R_{2}=1-\frac{1}{n}\right\}.$
@@
(2.11) $\mathcal{C}=\left\{\frac{1}{n-1}\left(G-H^{2}+d_{1}R_{1}+d_{2}R_{2}\right)=\frac{1-H^{2}}{n}\right\}.$
By a direct computation, one verifies that the algebraic surface $\mathcal{C}$ is invariant under the flow of (2.9), without appealing to (2.5). Recall that $Y$ and $Z$ encode metric components, so along the solutions of interest they are naturally non-negative. In the new coordinates, the range of the mean curvature is compressed from the entire real line to the open interval (1,1)(-1,1). These geometric constraints are reflected in the structure of the dynamical system, as stated below.
Proposition 2.1.
@@
is invariant under the flow of (2.9).
Proof.
Equation (2.9) shows that the sign of $Y$ and $Z$ is preserved, so we may restrict to $Y,Z\geq 0$. Moreover, differentiating $H$ gives
(2.12) $H^{\prime}=(H^{2}-1)\left(G+\frac{1}{n}(1-H^{2})\right)=(H^{2}-1)\left(\tfrac{1}{n}+\tfrac{d_{1}d_{2}}{n}(X_{1}-X_{2})^{2}\right).$
Hence the hypersurfaces $\{H=\pm 1\}$ are flow-invariant. In particular, if $H^{2}<1$ at some point along an integral curve, the integral curve cannot cross $\{H^{2}=1\}$, showing that $\mathcal{E}$ is flow-invariant. ∎
Remark 2.2.
The region $\{H^{2}<1\}$ corresponds to positive Einstein metrics ($\Lambda>0$). The system admits a $\mathbb{Z}_{2}$-symmetry
$(\eta,X_{1},X_{2},Y,Z)\longleftrightarrow(-\eta,-X_{1},-X_{2},Y,Z),$
which corresponds geometrically to reversing the normal direction of the hypersurface. This symmetry allows one to reflect an integral curve in $\mathcal{E}\cap\{H>0\}$ that intersects $\{X_{1}=X_{2}=0\}$ across $\{H=0\}$ and glue it to its mirror image, thereby producing a complete solution defined on both ends. For such a metric, the initial and terminal conditions of $f_{2}$ are identical, and we have $\underline{f}=\overline{f}$ in (2.7).
Remark 2.3.
In the Ricci-flat case ($\Lambda=0$), one may instead consider $d\eta=\mathrm{tr}(L)\,dt$ and redefine the variables $X_{1},X_{2},Y,Z$ accordingly. The resulting dynamical system coincides with the restriction of (2.9) to $\mathcal{RF}^{+}:=\mathcal{E}\cap\{H=1\}$. By the $\mathbb{Z}_{2}$-symmetry mentioned above, the subsystem restricted to $\mathcal{RF}^{-}:=\mathcal{E}\cap\{H=-1\}$ also corresponds to the Ricci-flat case after a time reversal.
To retrieve the original system, consider
(2.13) $t=\int_{-\infty}^{\eta}\sqrt{\frac{1-H^{2}}{n\Lambda}}d\tilde{\eta},\quad f_{1}^{2}=\frac{1-H^{2}}{n\Lambda Y^{2}},\quad f_{2}^{2}=\frac{1-H^{2}}{n\Lambda YZ}.$